OpenStudy (anonymous):

For what values of x are the following series convergent? In each case find the sum of the series for those values of x. a) 1 + (x-4) + (x-4)^2 + (x-4)^3 +... b) \[\sum_{n=1}^{\infty} 2^nx^n\]

OpenStudy (luigi0210):

What kind of test do you know to test for convergence?

OpenStudy (anonymous):

I'm not sure what test, but i know this formula, S= a/1-r.

OpenStudy (luigi0210):

Okay, haven't done this in a while.. sorry. But I assume you mean the Geometric Series of Convergence Test?

OpenStudy (luigi0210):

http://prntscr.com/657xlc

OpenStudy (anonymous):

yeah, it's geometric series

OpenStudy (luigi0210):

To test for the convergence just look at the value \(\color{red}{r}\)

OpenStudy (luigi0210):

If r is less than one, it converges. Greater than or equal to, it diverges.

OpenStudy (luigi0210):

I'm talking about b btw, not a :P

OpenStudy (anonymous):

I'm not sure how create the restrictions to make them converge.

OpenStudy (luigi0210):

I gtg, sorry :/ @ganeshie8 @Directrix @TheSmartOne @iambatman can help you :)

OpenStudy (anonymous):

Well, you can say: $$ \sum_{n=1}^\infty 2^n x^n = \sum_{n=1}^\infty (2x)^n = \sum_{n=1}^\infty u^n $$And now say: $$ u < 1 \implies 2x < 1 \implies x < \frac{1}{2} \\ \sum_{n=1}^\infty u^n = \frac{1}{1-u} = \frac{1}{1-2x} $$

OpenStudy (mathmath333):

@pitamar it should be \(\mid u \mid <1\) in your solution

OpenStudy (anonymous):

Yes I see it now, sorry =| @mathmath333

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{1 + (x-4) + (x-4)^2 + (x-4)^3 +\cdot \cdot \cdot \cdot \infty \hspace{.33em}\\~\\ (x-4)^0 + (x-4)^1 + (x-4)^2 + (x-4)^3 +\cdot \cdot \cdot \cdot \infty \hspace{.33em}\\~\\ \normalsize \text{let r=x-4} \hspace{.33em}\\~\\ (r)^0 + (r)^1 + (r)^2 + (r)^3 +\cdot \cdot \cdot \cdot \infty \hspace{.33em}\\~\\ \mid r \mid < 1 \hspace{.33em}\\~\\ \implies -1<r<1 \hspace{.33em}\\~\\ \implies -1<x-4<1 \hspace{.33em}\\~\\ \implies -1+4<x<1+4 \hspace{.33em}\\~\\ \implies 3<x<5 \hspace{.33em}\\~\\ }\end{align}\)